A quality control engineer at a potato chip company tests the bag filling machine by weighing bags of potato chips. Not every bag contains exactly the same weight. But if more than 11% of bags are over-filled then they stop production to fix the machine. They define over-filled to be more than 1 ounce above the weight on the package.

The engineer weighs 140 bags and finds that 37 of them are over-filled. He plans to test the hypotheses H0: p = 0.11 versus Ha: p > 0.11. What is the test statistic?

z =

(Please round your answer to two decimal places if necessary.)
Answer #1

This is the right tailed test .

The null and alternative hypothesis is

H0 : p =0.11

Ha : p > 0.11

\hat p = x / n = 37/140=0.264

P0 = 0.11

1 - P0 = 0.89

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.264-0.11/ [\sqrt(0.11*0.89) /140 ]

= 6.30